OpenStudy (abbycross167):

Can someone please help me to answer this math question? Please? I will fan and give a medal...

OpenStudy (abbycross167):

A laborer can do a certain job in 5 days, a second can do the job in 6 days, and a third in 8 days. In what time can the three together do the job?

OpenStudy (mertsj):

The first laborer does 1/5 jobs per day. In x days he does x/5 jobs.

OpenStudy (mertsj):

What does the second laborer do in one day?

OpenStudy (abbycross167):

1/6 = x/6? @Mertsj

OpenStudy (abbycross167):

@jim_thompson5910 would you mind helping me please?

jimthompson5910 (jim_thompson5910):

`A laborer can do a certain job in 5 days` the person can do 1/5 of the job in 1 day `a second can do the job in 6 days` the person can do 1/6 of the job in 1 day `a third in 8 days.` the person can do 1/8 of the job in 1 day agreed so far?

OpenStudy (abbycross167):

yes sir

jimthompson5910 (jim_thompson5910):

now add up all the fractions 1/5+1/6+1/8 = ???

OpenStudy (abbycross167):

59/120

jimthompson5910 (jim_thompson5910):

yes

jimthompson5910 (jim_thompson5910):

their combined rate is 59/120 jobs per day (combined rate)*(time) = 1 (59/120)*t = 1 solve for t

OpenStudy (abbycross167):

how would I do that?

jimthompson5910 (jim_thompson5910):

multiply both sides by the reciprocal of 59/120

jimthompson5910 (jim_thompson5910):

\[\Large \frac{59}{120}t = 1\] \[\Large \frac{120}{59}*\frac{59}{120}t = \frac{120}{59}*1\] \[\Large t = \frac{120}{59} \approx 2.0339\] so it will take them about 2.0339 days (a little over 2 whole days) if the three work together. This is assuming that they can work together and not get in each other's way, slow each other down, etc.

OpenStudy (abbycross167):

Thank you so much for the help!! You're amazing!

jimthompson5910 (jim_thompson5910):

np